OpenStudy (anonymous):

Which number equals sqrt ( 1,029 )

OpenStudy (anonymous):

Something around 32.

OpenStudy (ghazi):

32.07

OpenStudy (anonymous):

7 sqrt (21)

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

3 sqrt (21)

OpenStudy (anonymous):

21 sqrt (3)

OpenStudy (ghazi):

what are these?

OpenStudy (anonymous):

chhoices

OpenStudy (ghazi):

then break 1029

OpenStudy (anonymous):

1029=3*7*7*7 = 21*49, take out the square-root of 49 and 21 is left under the radical.

OpenStudy (anonymous):

7 sqrt 21

OpenStudy (anonymous):

Which expression is the simplified form of sqrt 567w^12